OpenStudy (fellowroot):

eliminating the parameter of the parametric equations. aka getting rid of t x = t^2 + t y = t^2 - t

jimthompson5910 (jim_thompson5910):

Add the two equations to get x = t^2 + t y = t^2 - t ----------- x+y = 2t^2+0t So we know that x+y = 2t^2

jimthompson5910 (jim_thompson5910):

solve x+y = 2t^2 for t and use this to plug it back into the two original equations (to eliminate t and replace it with an expression in terms of x and y)

OpenStudy (fellowroot):

But when you solve for t in x + y = 2t^2 you get t = +sqrt((x+y)/2) & t = -sqrt((x+y)/2) so which one do we plug in the + or -?

jimthompson5910 (jim_thompson5910):

well either one works, so usually the problem adds on something like "t is positive" or "t is nonnegative"

jimthompson5910 (jim_thompson5910):

I would go with the positive one since that's usually the one used anyway